# Revela: Dense Retriever Learning via Language Modeling

### v2

@@ 6 Future Directions
We train retrievers directly from raw text via self-supervised language modeling, sidestepping query-doc pairs, a breakthrough beyond the traditional paradigm. Based on this novel solution, We outline several future directions to suggest follow-up research. (1) Iterative Indexing: While Revela uses document chunking, a more general approach would iteratively index documents and group chunks by on-the-fly representations. Though explored in prior work (Izacard et al., 2023; Shi et al., 2024), the high computational cost of such methods remains a key challenge for future work. (2) Scaling up: We envision several directions to scale up Revela, including increasing retriever size in § 4.2, and increasing LM size in § 5. Additionally, incorporating advanced attention mechanisms (Yuan et al., 2025) may enhance retriever learning and accelerate the training. (3) Multi-modality: Although Revela targets text and code retrieval, this retriever-via-language-modeling paradigm can, in principle, generalize to other LM-modeled modalities, such as images (Jiang et al., 2025).
@@ Acknowledgments
We thank Kexin Wang, Jan Buchmann, Falko Helm, and Yuhao Zhang for the discussion and comments on an early draft of this work.
@@ B.4 Checkpoints
We include the off-the-shelf unsupervised and supervised retrievers and the corresponding huggingface URLs in Table 5.
@@ B.6 Supplementary Results
Table 7: Performance on BRIGHT (nDCG@10, %). Bold marks the best performance.
Dataset BM25 E5-PT Revela Revela REPLUG Revela E5-Mistral Revela Cohere OpenAI Voyage
Model Size 0.3B 0.5B 1B 3B 3B 7B 8B
Biology 18.9 18.9 17.0 17.1 7.1 25.3 18.6 24.4 18.7 23.3 23.1
Earth Science 27.2 29.2 28.8 29.1 13.1 40.0 26.0 40.9 28.4 26.7 25.4
Economics 14.9 10.3 13.9 13.2 12.9 17.1 15.5 21.7 20.4 19.5 19.9
Psychology 12.5 17.6 13.2 14.9 15.1 21.4 15.8 24.9 21.6 27.6 24.9
Robotics 13.6 7.0 9.1 11.0 7.6 15.5 16.3 14.6 16.3 12.8 10.8
StackOverflow 18.4 7.4 9.0 13.2 6.0 14.0 11.2 15.0 18.3 14.3 16.8
Sustainable 15.0 14.1 13.9 15.4 11.6 19.1 18.1 19.7 17.6 20.5 15.4
Pony 7.9 1.8 3.9 3.0 1.1 6.5 4.9 2.0 1.9 2.4 1.5
LeetCode 24.4 13.9 16.9 16.0 15.3 16.0 28.7 18.4 26.8 23.6 30.6
AoPS 6.2 1.4 4.6 2.7 5.6 3.8 7.1 3.5 6.3 8.5 7.5
ThmQA-Thm 4.9 4.9 5.6 7.3 8.5 9.7 26.8 11.8 7.2 11.7 11.6
ThmQA-Q 10.4 10.1 14.0 14.6 14.5 14.6 26.1 16.1 15.7 23.5 27.4
Mean 14.5 11.4 12.5 13.1 9.9 16.9 17.9 17.8 16.6 17.9 17.9
@@ C.2 Analysis on Batch sizes
Table 10: Retrieval performance (%, NDCG@10) across the 10 CoIR tasks. We report three encoder sizes (135M, 500M, 1B) and three batch sizes (bs4, bs8, bs16).
@@ C.3 Analysis on the Size of LMs
Table 13: Comparison of retrieval performance (%) across CoIR benchmark tasks for three LM model sizes. Bold in each row indicates the best performance in each task for one specific size of retriever.
@@ C.6 LM performance after Revela
We evaluate the original LLaMA-3.2-1B and the model co-trained with the retriever (LLaMA-3.2-1B backboned on Wikipedia in § 4.1) on seven tasks, including ARC (Clark et al., 2018), COPA (Roemmele et al., 2011), OpenBookQA (Mihaylov et al., 2018), PIQA (Mihaylov et al., 2018), RTE (Bisk et al., 2020), WiC (Dagan et al., 2005), and Winogrande (Pilehvar & Camacho-Collados, 2019). As shown in Tab. 18, LM’s performance has been greatly preserved.
@@ C.7 Comparison between Revelaand REPLUG
  • REPLUG: Computes pairwise relevance between all document pairs using LM perplexity, resulting in a forward cost of $B^{2}\times(2L)^{2}$, where the pair of texts is concatenated to pass through the LMs to calculate their relevance.
  • Revela: As detailed in § 3.4, uses in-batch attention with concatenated inputs, yielding a forward cost of $B\times(2L)^{2}$. If approximately taking backward cost as the same as forward, the cost is still linearly proportional to the batch size.

### v3

@@ 6 Future Directions
We train retrievers directly from raw text via self-supervised language modeling, sidestepping query-doc pairs, a breakthrough beyond the traditional paradigm. Based on this novel solution, We outline several future directions to suggest follow-up research. (1) Iterative Indexing: While Revela uses document chunking, a more general approach would iteratively index documents and group chunks by on-the-fly representations. Though explored in prior work (Izacard et al., 2023; Shi et al., 2024), the high computational cost of such methods remains a key challenge for future work. (2) Scaling up: We envision several directions to scale up Revela, including increasing retriever size in § 4.2, and increasing LM size in § 5. Additionally, incorporating more advanced attention mechanisms (Yuan et al., 2025) may enhance retriever learning and accelerate the training. (3) Multi-modality: Although Revela targets text and code retrieval, this retriever-via-language-modeling paradigm can, in principle, generalize to modalities, such as images (Jiang et al., 2025).
@@ Acknowledgments
We thank Kexin Wang, Sheng Lu, Jan Buchmann, Falko Helm, and Yuhao Zhang for the discussion and comments on an early draft of this work.
@@
This work has been supported by the European Union (ERC, InterText, 101054961). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The work is related to WP1 on cross-document modelling.
We gratefully acknowledge support from the hessian.AI Service Center (funded by the Federal Ministry of Research, Technology and Space, BMFTR, grant no. 16IS22091) and the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation, grant no. S-DIW04/0013/003).
@@ B.4 Checkpoints
We include the off-the-shelf unsupervised and supervised retrievers and the corresponding huggingface URLs in Tab. 5. For the prosperity APIs, we list the URLs in Tab. 6.
@@
Table 6: Embedding models and their reference URLs.
Name URL
OpenAI-Ada-002 platform.openai.com/docs/guides/embeddings
Voyage-code-2 blog.voyageai.com/2024/01/23/voyage-code-2-elevate-your-code-retrieval/
text-embedding-3-large openai.com/index/new-embedding-models-and-api-updates/
voyage-large-2-instruct docs.voyageai.com/docs/embeddings
cohere-embed-english-v3.0 huggingface.co/Cohere/Cohere-embed-english-light-v3.0
@@ B.6 Supplementary Results
Table 8: Performance on BRIGHT (nDCG@10, %). Bold marks the best performance. The results of BM25, E5-Mistral and APIs are taken from BRIGHT (Hongjin et al., 2025).
Dataset BM25 E5-PT Revela Revela REPLUG Revela E5-Mistral Cohere OpenAI Voyage
Model Size 0.3B 0.5B 1B 3B 3B 7B
Biology 18.9 18.5 16.9 16.7 6.8 24.9 18.6 18.7 23.3 23.1
Earth Science 27.2 29.5 29.0 29.4 13.0 40.3 26.0 28.4 26.7 25.4
Economics 14.9 10.2 13.9 13.2 12.9 17.2 15.5 20.4 19.5 19.9
Psychology 12.5 17.6 13.3 15.0 15.1 21.4 15.8 21.6 27.6 24.9
Robotics 13.6 7.1 9.1 10.5 7.4 15.3 16.3 16.3 12.8 10.8
StackOverflow 18.4 8.7 10.8 15.5 7.5 16.9 11.2 18.3 14.3 16.8
Sustainable 15.0 14.2 13.9 15.4 12.1 19.2 18.1 17.6 20.5 15.4
Pony 7.9 1.8 3.9 3.0 1.1 6.5 4.9 1.9 2.4 1.5
LeetCode 24.4 22.5 27.6 26.1 24.8 26.6 28.7 26.8 23.6 30.6
AoPS 6.2 3.3 12.5 8.8 13.3 11.6 7.1 6.3 8.5 7.5
ThmQA-Thm 4.9 6.0 6.7 8.9 9.2 14.2 26.8 7.2 11.7 11.6
ThmQA-Q 10.4 17.4 24.8 24.1 22.9 27.1 26.1 15.7 23.5 27.4
Mean 14.5 13.1 15.2 15.6 12.2 20.1 17.9 16.6 17.9 17.9
@@ C.2 Analysis on Batch sizes
Table 11: Retrieval performance (nDCG@10, %) across the 10 CoIR tasks. We report three encoder sizes (135M, 500M, 1B) and three batch sizes (bs4, bs8, bs16).
@@ C.3 Analysis on the Size of LMs
Table 14: nDCG@10 (%) on CoIR for three model sizes (Revela0.1B, Revela0.5B, Revela1B), with LM sizes in ascending order. Bold in each row indicates the best performance for each retriever size.
@@ C.6 LM performance after Revela
We evaluate the original LLaMA-3.2-1B and the model co-trained with the retriever (LLaMA-3.2-1B backboned on Wikipedia in § 4.1) on seven tasks, including ARC (Clark et al., 2018), COPA (Roemmele et al., 2011), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), RTE (Dagan et al., 2005), WiC (Pilehvar & Camacho-Collados, 2019), and Winogrande (Sakaguchi et al., 2020). As shown in Tab. 19, LM’s performance has been greatly preserved.
@@ C.7 Comparison between Revela and REPLUG
  • REPLUG: To compute pairwise relevance between all document pairs using LM perplexity requires processing each pair independently. Each pair is concatenated (length $2L$) and passed through the LM, resulting in a forward cost of $\mathcal{O}(B^{2}\times(2L)^{2})$ for all B(B1)/2B(B-1)/2 pairs.
  • Revela: As detailed in § 3.4, uses in-batch attention with concatenated inputs. By processing all $B$ documents jointly in a single forward pass with cross-attention, it computes inter-document relevance with a cost of $\mathcal{O}(B\times(2L)^{2})$. Including backward passes (which scale similarly), the total cost remains linear in batch size.